# Low-Complexity OTFS-Based Over-the-Air Computation Design for Time-Varying Channels

### v1

@@ VI-A Error Performance for S1, S2, and S3
In Fig. 4, we plot MSE versus SNR for S1 and S2, where SNR is defined as $P_{\mathrm{s}}/\sigma^{2}$ and $P_{\mathrm{s}}$ is the average power budget. First, we apply different transmission power policies for S1. Our observations reveal that assigning optimal transmit power according to (25) results in a significantly lower MSE compared to allotting full transmit power to all devices or to a singular user. This supports the optimal efficiency of $p_{u}^{*}$ and $\eta^{*}$ in minimizing MSE. Second, we observe that an MSE plateaus of S1 for larger SNR, which reveals the fact that interference plays the pivotal role in degrading the performance of estimation. We further observe that MSE for S2 is much lower than that of S1, which proves the superiority of ZP-assisted OTFS in mitigating the impact of interference. We also observe that this superiority becomes more obvious as SNR increases. We note that this superiority is established on the premise of sacrificing communication efficiency.
@@
In Fig. 5, we compare the MSE for our proposed estimation methods with benchmarks. First, Fig. 5 shows that our proposed estimation schemes S1 and S2 have a MSE superiority compared to the methods in [17, 18], and this MSE superiority of S1 only holds for low SNR. Second, we observe that the MSE performance for our proposed methods is always better than that when only considering the precoding matrix at each device or only considering the receiving filter at the FC. This observation illustrates the necessity of considering both the pre-processing and post-processing at the TX and RX.
@@
Figure 6: MSE versus number of devices for S1 and S2, where each path has a different delay or two paths have the same delay and $\mathrm{SNR}=10\;\mathrm{dB}$.
@@
In Fig. 6, we plot the MSE versus the number of devices participating in AirComp. First, we observe a reduction in MSE as the number of devices increases, aligning with the general understanding that increased device participation enhances estimation accuracy in AirComp. Second, we consider the scenario where two paths share the identical delay. We first assume $l_{1}=l_{2}\neq l_{3}\neq l_{4}$, which leads to two non-zero values in each row of $\boldsymbol{K}_{m,l_{1}}^{u}$. As $\boldsymbol{K}_{m,l_{1}}^{u}$ directly multiplies with the estimated function $\sum_{u=1}^{U}\boldsymbol{x}_{u,m}$, there exists interference that cannot be canceled during estimation, causing a slight rise in MSE compared to scenarios with varying delays in each path. We then assume $l_{2}=l_{3}\neq l_{1}\neq l_{4}$, which leads to two non-zero values in each row of $\boldsymbol{K}_{m,l_{2}}^{u}$. As $\boldsymbol{K}_{m,l_{2}}^{u}$ multiplies only with previously estimated interferences, effective interference cancellation occurs during the current estimation process. Moreover, the sharing of identical values by two paths enlarges the average pairwise differences between $l_{1}$, $l_{2}$, and $l_{4}$. Consequently, these two factors contribute to a lower MSE for $l_{2}=l_{3}$ compared to scenarios where each path experiences varying delays.
Figure 7: MSE versus number of paths for S1 and S2, where $\mathrm{SNR}=10\;\mathrm{dB}$.
@@
In Fig. 7, we plot MSE versus the number of paths for S1 and S2. For S2, we consider the scenarios that either each path has a different delay or two paths share the same delay. For different delay in each path, the number of paths is from 1 to 11. for two paths sharing the same delay, the number of paths is from 3 to 11. First, we observe that the MSEs for S1 and S2 increase as the number of paths increases. This is due to the fact that a larger number of paths introduces more interference terms. Second, the MSE gap between S1 and S2 decreases as the number of paths increases, which indicates that the SIC-based algorithm in S2 is more proficient in mitigating the interference when the number of channel paths is small. Third, for the scenario that two paths share the same delay, we consider either $l_{1}=l_{2}$ or $l_{2}=l_{3}$. We observe that the MSE for $l_{1}=l_{2}$ is always larger than that for each path having a different delay. Eventually, when the number of paths is sufficiently large, the MSE performance is the same for scenarios where $l_{1}=l_{2}$, $l_{2}=l_{3}$, and each path has a different delay.
@@
In Fig. 10, we evaluate the error performance of S3 for different number of channel paths and the moving speed of the devices, denoted by $v_{u}$. First, we observe that the MSE increases with the increase in the moving speed of devices. This is due to the fact that a higher moving speed leads to a higher Doppler shift which degrades the estimation performance at FC. Second, when $\mathrm{SNR}=20\;\mathrm{dB}$, we observe a more significant increase in MSE when the velocity $v_{u}$ rises from $25\;\mathrm{km}/\mathrm{h}$ to $127\;\mathrm{km}/\mathrm{h}$ compared to the increase from $127\;\mathrm{km}/\mathrm{h}$ to $253\;\mathrm{km}/\mathrm{h}$. This observation suggests that OTFS demonstrates greater robustness to velocity changes in the high-speed range. Third, we observe that the MSE improves with the increase in the number of channel paths between each device and FC. This is because a larger number of paths introduces more ISI and ILI.
@@ VII Conclusion
In this paper, we considered an OTFS-based AirComp over double-selective time-varying channels, where mobile devices first modulate their data on the delay-Doppler (DD) domain and then transmit them to the fusion center (FC) for averaging. In particular, we proposed three schemes S1, S2, and S3 of increasing complexity for the target function estimation. In S1, we derived the closed-form expressions for the optimal transmit power and denoising factor. In S2, we applied a ZP-assisted OTFS-based AirComp, where we proposed an SIC-based algorithm and derived the closed-form expressions for the optimal transmit power, denoising factor, and the weight for interference cancellation. In S3, we proposed an iterative algorithm to jointly optimize the precoding matrix and receiving filter matrix. Numerical results validated the superiority of the proposed schemes in the error performance than benchmarks. They also showed a low computational complexity of S1 and S2 compared to benchmarks. Additionally, given the shared parameters between OTFS and radar sensing, such as delay and Doppler, the application of OTFS in Integrated Sensing and Communication (ISAC) has garnered considerable interest. Building on the results of this paper, an intriguing area of investigation emerges in exploring an integrated approach to sensing, communication, and computation within the DD domain.
@@ Appendix B Proof of Theorem 2
We note that the optimization problem in (B) is the same as that in [9]. Then, (41) and (42) can be obtained following [9, Eqs. (24), (25)], whose derivation processes are omitted here due to the page limit.

### v2

@@ VI-A Error Performance for S1, S2, and S3
In Fig. 4, we plot MSE versus SNR for S1 and S2, where SNR is defined as $P_{\mathrm{s}}/\sigma^{2}$ and $P_{\mathrm{s}}$ is the average power budget. First, we apply different transmission power policies for S1. Our observations reveal that assigning optimal transmit power according to (26) results in a significantly lower MSE compared to allotting full transmit power to all devices or to a singular user. This supports the optimal efficiency of $p_{u}^{*}$ and $\eta^{*}$ in minimizing MSE. Second, we observe that an MSE plateaus of S1 for larger SNR, which reveals the fact that interference plays the pivotal role in degrading the performance of estimation. We further observe that MSE for S2 is much lower than that of S1, which proves the superiority of ZP-assisted OTFS in mitigating the impact of interference. We also observe that this superiority becomes more obvious as SNR increases. We note that this superiority is established on the premise of sacrificing communication efficiency. Specifically, the reduction in transmission rate (measured in symbols per second) for ZP-assisted OTFS is given by $l_{u,\mathrm{max}}/T$.
@@
In Fig. 5, we compare the MSE for our proposed estimation methods with benchmarks. First, Fig. 5 shows that our proposed estimation schemes S1 and S2 have a MSE superiority compared to the methods in [18, 19], and this MSE superiority of S1 only holds for low SNR. Second, we observe that the MSE performance for our proposed methods is always better than that when only considering the precoding matrix at each device or only considering the receiving filter at the FC. This observation illustrates the necessity of considering both the pre-processing and post-processing at the TX and RX. Third, the MSEs for methods that consider only precoding or only the receiving filter remain nearly constant as the SNR changes. This occurs because using either precoding or a receiving filter alone fails to adequately compensate for signal distortion caused by channel fading, delay, and Doppler effects, resulting in errors significantly larger than those due to channel noise. This observation underscores the critical importance of incorporating both precoding and receiving filters to effectively recover the target function.
@@
Figure 6: MSE versus number of devices, S1, S2, and the methods in [18, 19]. For S2, each path has a different delay or two paths have the same delay and $\mathrm{SNR}=10\;\mathrm{dB}$.
@@
In Fig. 6, we plot the MSE versus the number of devices participating in AirComp. First, we observe a reduction in MSE as the number of devices increases, aligning with the general understanding that increased device participation enhances estimation accuracy in AirComp. Second, we consider the scenario where two paths share the identical delay. We first assume $l_{1}=l_{2}\neq l_{3}\neq l_{4}$, which leads to two non-zero values in each row of $\boldsymbol{K}_{m,l_{1}}^{u}$. As $\boldsymbol{K}_{m,l_{1}}^{u}$ directly multiplies with the estimated function $\sum_{u=1}^{U}\boldsymbol{x}_{u,m}$, there exists interference that cannot be canceled during estimation, causing a slight rise in MSE compared to scenarios with varying delays in each path. We then assume $l_{2}=l_{3}\neq l_{1}\neq l_{4}$, which leads to two non-zero values in each row of $\boldsymbol{K}_{m,l_{2}}^{u}$. As $\boldsymbol{K}_{m,l_{2}}^{u}$ multiplies only with previously estimated interferences, effective interference cancellation occurs during the current estimation process. Moreover, the sharing of identical values by two paths enlarges the average pairwise differences between $l_{1}$, $l_{2}$, and $l_{4}$. Consequently, these two factors contribute to a lower MSE for $l_{2}=l_{3}$ compared to scenarios where each path experiences varying delays. Third, comparing the methods in [18, 19] with S1 and S2 reveals that the methods in [18, 19] outperform S1 when the number of devices is large, while S2 consistently achieves a lower MSE than the benchmarks in [18, 19].
Figure 7: MSE versus number of paths for S1, S2, and methods in [18, 19], where $\mathrm{SNR}=10\;\mathrm{dB}$.
@@
In Fig. 7, we plot MSE versus the number of paths for S1, S2, and benchmarks in [18, 19]. For S2, we consider the scenarios that either each path has a different delay or two paths share the same delay. For different delay in each path, the number of paths is from 1 to 11. for two paths sharing the same delay, the number of paths is from 3 to 11. First, we observe that the MSEs for S1 and S2 increase as the number of paths increases. This is due to the fact that a larger number of paths introduces more interference terms. Second, the MSE gap between S1 and S2 decreases as the number of paths increases, which indicates that the SIC-based algorithm in S2 is more proficient in mitigating the interference when the number of channel paths is small. Third, for the scenario that two paths share the same delay, we consider either $l_{1}=l_{2}$ or $l_{2}=l_{3}$. We observe that the MSE for $l_{1}=l_{2}$ is always larger than that for each path having a different delay. Eventually, when the number of paths is sufficiently large, the MSE performance is the same for scenarios where $l_{1}=l_{2}$, $l_{2}=l_{3}$, and each path has a different delay. Fourth, the MSEs of S1 and S2 are lower than those of [18, 19] when the number of paths exceeds 3, demonstrating the superiority of our proposed methods in handling a large number of channel paths.
@@
In Fig. 10, we evaluate the error performance of S3 for different number of channel paths and the moving speed of the devices. For each $v_{u}$, we calculate $k_{\mathrm{max}}$ as $k_{\mathrm{max}}=v_{u}f_{\mathrm{c}}N/(c\Delta f)$. Then, the Doppler index of each path is randomly generated within the range $[-k_{\mathrm{max}},k_{\mathrm{max}}]$. First, we observe that the MSE increases with the increase in the moving speed of devices. This is due to the fact that a higher moving speed leads to a higher Doppler shift which degrades the estimation performance at FC. Second, when $\mathrm{SNR}=20\;\mathrm{dB}$, we observe a more significant increase in MSE when the velocity $v_{u}$ rises from $25\;\mathrm{km}/\mathrm{h}$ to $127\;\mathrm{km}/\mathrm{h}$ compared to the increase from $127\;\mathrm{km}/\mathrm{h}$ to $253\;\mathrm{km}/\mathrm{h}$. This observation suggests that OTFS demonstrates greater robustness to velocity changes in the high-speed range. Third, we observe that the MSE improves with the increase in the number of channel paths between each device and FC. This is because a larger number of paths introduces more ISI and ILI.
@@ VI-B Computational Complexity
In summary, each scheme offers unique advantages. S1 excels in simplicity, low computational complexity, and robustness to varying OTFS size and the number of channel paths. S2 strikes a balance with better error performance than S1 and lower complexity than S3. S3 offers the best error performance but at the cost of higher computational complexity.Ultimately, there is a trade-off between computational complexity and error performance, and the choice of scheme depends on the specific requirements of the scenario.
@@ VII Conclusion
In this paper, we considered an OTFS-based AirComp over double-selective time-varying channels, where mobile devices first modulate their data on the delay-Doppler (DD) domain and then transmit them to the fusion center (FC) for averaging. In particular, we proposed three schemes S1, S2, and S3 of increasing complexity for the target function estimation. In S1, we derived the closed-form expressions for the optimal transmit power and denoising factor. In S2, we applied a ZP-assisted OTFS-based AirComp, where we proposed an SIC-based algorithm and derived the closed-form expressions for the optimal transmit power, denoising factor, and the weight for interference cancellation. In S3, we proposed an iterative algorithm to jointly optimize the precoding matrix and receiving filter matrix. Numerical results validated the superiority of the proposed schemes in the error performance than benchmarks. They also showed a low computational complexity of S1 and S2 compared to benchmarks. Future research directions include conducting real-world OTFS measurements in vehicular communication scenarios to evaluate the practical implementation of the proposed algorithms. Additionally, given the shared parameters between OTFS and radar sensing, such as delay and Doppler, the application of OTFS in Integrated Sensing and Communication (ISAC) has garnered considerable interest. Building on the results of this paper, an intriguing area of investigation emerges in exploring an integrated approach to sensing, communication, and computation within the DD domain.
@@ Appendix B Proof of Theorem 2
We note that the optimization problem in (B) is the same as that in [10]. Then, (42) and (43) can be obtained following [10, Eqs. (24), (25)], whose derivation processes are omitted here due to the page limit.